Grande03:

In a experiment, it was found that the population of a bacterial colony doubles every 10 minutes. The following function represents the number of bacteria in the colony after x minutes. 3(2)^x/10 Which statements are true? - The expression (1,024)^x reveals the approximate rate of increase in the population of the bacterial colony per hour. - The expression (0.93)^x reveals the approximate rate of increase in the population of the bacterial colony per minute. - The expression (1.07)^x reveals the approximate rate of increase in the population of the bacterial colony per minute. - The expression 2^x/10 reveals that the population of the bacterial colony increases by 100% every 10 minutes. - The expression (1,024)^x reveals the approximate rate of increase in the population of the bacterial colony per minute. - The expression 2^x/10 reveals that the population of the bacterial colony increases by 200% every 10 minutes.

surjithayer:

what do you think solution can be?

surjithayer:

let us check \[3(2)^{\frac{x}{10}}\] when x=0.\[population ~of ~bacteria~=3(2)^0=3\] when x=10\[population=3(2)^{\frac{x}{10}}=3(2)^{\frac{10}{10}}=3(2)^1=6\] increase in population=6-3=3 \[percentage~increase= \frac{ increase~of~poulation }{ original~population} \times ~100\] plug the values and find the solution.